# MKF-ADS: A Multi-Knowledge Fused Anomaly Detection System for Automotive

### 07 Mar 2024 v1

@@ III-B Attack Model
A factual belief that an attacker can remotely compromise one or more ECUs via a wireless interface (e.g., a telemetries port  [52]) or physically (e.g., via OBD-II  [53]). Therefore, every ECU may become the potential entrance for the attacker to access the CAN bus and generate various attacks. The details of these attack models are summarized as follows:
  • Denial of Service (DoS) Attack: The attack aims to refuse or interfere with the intended function of the CAN bus. The attacker usually sends a large number of redundant messages with higher priority to exhaust system resources and paralyze the functions. Also, there is an insidious operation that mimics a specific CAN ID to transmit the same CAN messages, with a relatively increasing frequency rate.
@@ IV-C2 PatchST Module
Transform-based high-level context knowledge is a state-of-the-art approach for anomaly detection in multivariate temporal sequences, due to its long-range modeling capability. Inspired by [55], we introduce PatchST and regard it as complementary to spatial-temporal knowledge, relaxing the restriction that makes it difficult to use complex models for resource-constrained IVNs [56, 57, 23].
@@ V-C Generalization performance
Figure 13: Evaluation of the ROAD dataset, including four spoofing attacks and four masquerade attacks (related to the Engine, Speedometer, and Light), where the red line is the ground truth, and the green line is MKF-ADS prediction.
@@ VI Conclusion and Future Directions
In this paper, we propose a novel multi-knowledge fused anomaly detection model, which can efficiently predict CAN messages that deviate from normal behavior. This model is developed to integrate spatial-temporal correlation with an attention mechanism (STcAM) module and patch sparse-transformer module (PatchST), achieving a knowledge-fused expression on the representation level. The STcAM with fine-pruning uses Conv1D to extract spatial features and subsequently utilizes the Bi-LSTM to obtain the temporal features, of which the attention mechanism will focus on the important time steps. Meanwhile, this work exploits PatchST to capture the combined long-time historical features from independent univariate time series. The framework is based on knowledge distillation to STcAM as a student model for learning intrinsic knowledge and cross the ability to mimic PatchST. As only deploying STcAM, the proposed model is adapted to a resource-limited IVN environment. Compared with the baseline in this paradigm, the prediction error and detection capability are improved and robust.

### 15 Mar 2024 v2

@@ III-B Attack Model
In this study, we consider that an attacker can remotely compromise one or more ECUs via a wireless interface (e.g., a telemetries port  [37]) or physically (e.g., via OBD-II  [53]). Therefore, every ECU may become the potential entrance for the attacker to access the CAN bus and generate various attacks. The details of these attack models are summarized as follows:
  • Denial of Service (DoS) Attack: The attack aims to refuse or interfere with the intended function of the CAN bus. The attacker usually sends a large number of redundant messages with higher priority to exhaust system resources and paralyze the functions.
@@ IV-C2 PatchST Module
Transform-based high-level context knowledge is a state-of-the-art approach for anomaly detection in multivariate temporal sequences, due to its long-range modeling capability. Inspired by [55], we introduce PatchST and regard it as complementary to spatial-temporal knowledge, relaxing the restriction that makes it difficult to use complex models for resource-constrained IVNs [26, 56].
@@ V-C Generalization performance
Figure 13: Evaluation of the ROAD dataset, including four spoofing attacks and four masquerade attacks (related to the Engine, Speedometer, and Light), where the red line is the ground truth, and the green line is prediction of MKF-ADS.
@@ VI Conclusion and Future Directions
In this paper, we propose a multi-knowledge fusion-based self-supervised anomaly detection model, which can efficiently predict CAN messages that deviate from normal behavior. This model is developed to integrate spatial-temporal correlation with an attention mechanism (STcAM) module and patch sparse-transformer module (PatchST), achieving a knowledge-fused expression on the representation level. The STcAM with fine-pruning uses Conv1D to extract spatial features and subsequently utilizes the Bi-LSTM to obtain the temporal features, of which the attention mechanism will focus on the important time steps. Meanwhile, this work exploits PatchST to capture the combined long-time historical features from independent univariate time series. The framework is based on knowledge distillation to STcAM as a student model for learning intrinsic knowledge and cross the ability to mimic PatchST. As only deploying STcAM, the proposed model is adapted to a resource-limited IVN environment. Compared with the baseline in this paradigm, the prediction error and detection capability are improved and robust.